Find an annual report for a company of interest (perhaps your employer). Publicly held company annual reports (10K) can be located via Edgar Archives on the Security and Exchange website.

To perform the search in the Edgar Archives you can use either the company name or "ticker symbol" for the company. [View instructions on how to access the annual reports]

Click on the link to "Interactive Data" and then the link to the "Financial Statements". Locate and review the Income Statement and Balance Sheet. Title your post with the name of your company and share the following in your post:

  1. What is the name of your company?
  2. What is the web address that leads to your financial reports?
  3. Why are you interested in this company?
  4. Calculate the gross profit percentage.
  5. Calculate the inventory turnover & days’ sales in inventory ratio.
  6. What have you learned from this analysis?
  • After you have conducted your research, please share with the class approximately 1-2 paragraphs.
Answer #1
What is the name of your company? MICROSOFT CORPORATION
What is the web address that leads to your financial reports? Visit microsoft's website, open investor section, go to SEC filings
Why are you interested in this company? Microsoft has always been an innovator. As of January 2019, more than 75% of active PCs use Windows as their operating system. Also Microsoft has made amazing advances in the field of computer gaming and who can forget the X-box? Microsoft always tries to stay ahead. I also admire microsoft and Mr. Gates for the philanthropic work undertaken. They have touched so many lives directly and indirectly.
Calculate the gross profit percentage. = Gross Margin / Total Revenue   = 72,007 / 110,360 = 65.25%
Calculate the inventory turnover & days’ sales in inventory ratio. Assuming all sales to be credit sales, Inventory turnover ratio is

= Credit Sales / (Opening Inventory + Closing Inventory)/2
= 110,360 / (2,181 + 2,662)/2
= 45.58 times

Days sales in inventory = 1/45.58 X 365 = 8.01 days
What have you learned from this analysis? This analysis tells us that microsoft has extremely good inventory turnover ratio. It takes an amazingly fast a little over 8 days to turn its inventory into sales. It is exceptionally well.